Use cylindrical shells to find the volume of the solid that results when the red region is revolved about theΒ y

y-axis. f(x)=x^2 ,a=1 ,b=4


1
Expert's answer
2021-10-11T15:50:35-0400

"V=\int_a^b 2\pi x f(x) dx\\
V=2\pi\int_1^4x^3dx\\
V=2\pi\left[\frac{x^4}{4}\right]_1^4\\
V=2\pi(64-\frac{1}{4})\\
V=\frac{255}{2}\pi"